The width of a rectangular play ground is 2x - 5 feet, and the length is 3x + 9 feet. Write a polynomial P(x) that represents the perimeter and then evaluate this perimeter polynomial if x is 4 feet.
Answer by checkley77(12844)

W=(2X+5)
L=(3X+9)
2L+2W=P
2(3X+9)+2(2X+5)=P
2(3*4+9)+2(2*4+5)=P
2(12+9)+2(8+5)=P
2*21+2*13=P
42+26=P
P=68 FT. ANS.